integrate x2 +x2 +5/(x2+4) (x+1) dx
1
Expert's answer
2020-10-07T15:03:39-0400

I = "\int" (x2 + x2 + "\frac{5}{(x^2+4)(x+1)}" )dx

Converting in partial fractions form, we get-

"\frac{5}{(x^2+4)(x+1)} = \frac{A}{(x+1)}+\frac{Bx+C}{(x^2+4)}"

5 = (A+B)"x^2" + (B+C)x + (4A+C)

Equating the coefficients and solving equations, we get

A = 1 ; B = -1 ; C = 1

I = "\int" (x2 + x2 + "\frac{1}{(x+1)}-\frac{x-1}{x^2+4}" )dx

I = "\int" (2x2 + "\frac{1}{(x+1)}-\frac{x}{(x^2+4)}+\frac{1}{(x^2+4)}" )dx

Since, integration of "\frac{1}{x+1}" = ln(x+1) , integration of "\frac{x}{x^2+4} is \frac{1}{2}ln(x^2+4)" and integration of "\frac{1}{x^2+4}" is "\frac{1}{2}tan^{-1}\frac{x}{2}" , we have

"I = \frac{2x^3}{3}+ln(x+1)-\frac{1}{2}[ln(x^2+4)-tan^{-1}(\frac{x}{2})]+const"